OpenStudy (anonymous):

find an equation of a line passing through the origin and parallel to the line passing through the points (2,8), (4,12)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(2\quad ,&8)\quad &(4\quad ,&12) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

i need a tutor this time around write the equation 5x-6y-12=0 in the slope intercept form and then find the slope and y intercept of the corresponding line